OpenStudy (anonymous):

What is the range of the graph of y = –2(x – 3)2 + 7?

OpenStudy (anonymous):

Graph the equation and determine the range of y values

OpenStudy (haseeb96):

y=-2(x^2 -6x +9) +7 y=-2x^2 -12x -18 +7 y= -2x^2 -12x -11 Its range is All real Numbers